OpenStudy (anonymous):

sq root of 2 X sq rt of 15 = how do you go about when trying to solve these type of problems i know sq rt of 64 more simple but i find this more complex

OpenStudy (dumbcow):

use this property \[\sqrt{a}*\sqrt{b} = \sqrt{a*b}\]

OpenStudy (anonymous):

ok this would give me an answer of sq rt 30 am i doing this correctly?

OpenStudy (dumbcow):

yes